OpenStudy (anonymous):

A monument is built out of toothpicks. There are 30 rows and each row has seven more toothpicks than the row above. The top row has five toothpicks. Part 1: How many toothpicks are on the bottom row? Part 2: Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (johnweldon1993):

Well not counting the top row....there are 29 rows.... adding 7 each time....so it would be 7 times 29 = 203 toothpicks I believe

OpenStudy (jdoe0001):

heheh, that'd be a small building, is not a good idea to make them off toothpicks btw :) 7 * 29 is 203 but the top row is 5 5+7 5+7+7 5+7+7+7 .....

OpenStudy (anonymous):

im thinkin 208

OpenStudy (johnweldon1993):

Or you can look at the equation \[a _{n}=a _{1}+(n-1)d\] 'd' is the common difference (7) n = number of rows you have (30) and a1 is your first term so \[a _{30}=5+(30-1)7\] Right @physicsboffin

OpenStudy (anonymous):

LOl @johnweldon1993

OpenStudy (johnweldon1993):

Which does come out to 208

OpenStudy (anonymous):

a+29d=5 d=-7 a=29(7) +5 =208

OpenStudy (cwrw238):

the formula for the nth term is a + (n - 1)d where a = first term, d = common difference

OpenStudy (cwrw238):

so the bottom term which is 30th is 5 + (30 - 1)*7 = 5 + 203 = 208